OpenStudy (anonymous):

Marie sells magazine subscriptions. She earns $15 a week plus $3 for each subscription she sells. How many subscriptions must she sell to earn $90 in one week?

OpenStudy (anonymous):

lol well the answer is 25

OpenStudy (anonymous):

well how did you get that answer?

OpenStudy (anonymous):

90 - 15= 75 75/3= 25

OpenStudy (anonymous):

all right thanks

OpenStudy (anonymous):

Solve the following for x, th number of magazines.\[3x+15=90\]

OpenStudy (anonymous):

ur welcome